Calculate the volume of Xe which could be isolated from 1000km3 sample air at standard temperature and pressure conditions,if the amount of Xe in the atmosphere is 0.076? Please explain with reasons..

Grade:12

1 Answers

Arun
4 years ago
Dear student

Volume v of air is 1000km3=1*10^12 m3.

Standard temperature, T=273.15K

Standard pressure P is 101325Pa.

0.076g of Xe will have (0.076/131.29) moles=5.789*10^-5.

The volume of 0.076g of xe is  7.6*10^-5/5.761=1.319*10^-5

Volume of Xe in 1000km3 is given by;

P/TRn=VV=nRT/P

V= (5.789*10^-4) *8.314* 273.15/101325

v =1.297*10^-5m3

Hope it helps